OpenStudy (anonymous):

Solve the system y=2.6x−3.4 −1.6x+0.4y=−0.8

OpenStudy (anonymous):

well it is easier to use substitution in this occasion? do you know how?

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

do i just plug the y equation (1st equation into the second?

OpenStudy (anonymous):



OpenStudy (anonymous):

can you solve now?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

THANK U SOO MUCH

OpenStudy (anonymous):

ur welcome :)

OpenStudy (anonymous):

DID U SOLVE THIS ONE?

OpenStudy (anonymous):

I WAN TO COMPARE ANSWERS

OpenStudy (anonymous):

ok so i got x=-1

OpenStudy (anonymous):

and y=-6

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

I GOT DECIMALS. (0.8, -1.32)

OpenStudy (anonymous):

really how?

OpenStudy (anonymous):

ILL SHOW YOU 1.6X+0.4(2.6X-3.4)=0.8 1.6+1.04X-1.36=0.8 2.64X-1.36=0.8 +1.36 +1.36 2.64X=2.16 2.64X= 2.64 X= 0.81

OpenStudy (anonymous):

and these you got were wrong, right?

OpenStudy (anonymous):

were mine wrong?

OpenStudy (anonymous):

YOURS WERE CORRECT

OpenStudy (anonymous):

I DONT KNOW WHAT IM DOING WRONG

OpenStudy (anonymous):

ok wait up, let me check :)

OpenStudy (anonymous):

KK

OpenStudy (anonymous):

oh i see , so this is ur equation right? 1.6X+0.4(2.6X-3.4)=0.8

OpenStudy (anonymous):

ur setup ^^^ right?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

YES. -0.8

OpenStudy (anonymous):

i believe you missed the -0.8 at the end after the = you put = 0.8 but its supposed to be = -0.8

OpenStudy (anonymous):

0K

OpenStudy (anonymous):

can you solve and then tell me what you get then?

OpenStudy (anonymous):

0.56/2.64=0.21 :(

OpenStudy (anonymous):

I DONT GET A NEGATIVE AT ALL

OpenStudy (anonymous):

ok sorry but its supposed to be: -1.6X+0.4(2.6X-3.4)=-0.8 you just made a simple mistake of missing the negative in front of -1.6 and -0.8

OpenStudy (anonymous):

ok can you solve now: -1.6X+0.4(2.6X-3.4)=-0.8 try it i guarantee you, that you will get the correct answer now :)

OpenStudy (anonymous):

OK:)

OpenStudy (anonymous):

I GOT -6

OpenStudy (anonymous):

THANK YOU!!!!!! FOR BEING SO PATIENT

OpenStudy (anonymous):

The Bread Alone Bakery has a daily overhead of $90. It costs $0.60 to bake each loaf of bread, and the bread sells for $1.50 per loaf. a. Write an equation for the cost, C, in terms of the number of loaves, x. b. write an equation for the revenue R

OpenStudy (anonymous):

can u help me with this one ?